OpenStudy (anonymous):

pleeeease help, will medal and fan(: Describe the transformations of the function f(x) = 3/(x+2) +4

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Do you how to graph a function?

OpenStudy (anonymous):

yes i know how to graph it, no i dont have any ideas. frankly, i dont quite understand what the question is asking`

OpenStudy (skullpatrol):

Compare the given function to the graph of f(x) = 1/x.

OpenStudy (skullpatrol):

And describe the changes :)

OpenStudy (anonymous):

i mean, i get that its 3 instead of 1 divided by x+2 instead of x and it adds 3, but how does that affect the graph?

OpenStudy (anonymous):

*adds 4

OpenStudy (skullpatrol):

"How does that effect the graph" is exactly what the question is asking for. First graph F(x) = 1/x. Then graph f(x) and see for yourself the differences :)

OpenStudy (anonymous):

hey, we have the same path of thinking haha. i did, but i dont know how to describe it.

OpenStudy (anonymous):

i mean, it moves up and to the left but im assuming my teacher wants more than that. how do i say how many units its moved up when it has no definitive vertex?

OpenStudy (skullpatrol):

Use phrases like "shift up, down, left, or right" and "compress or stretch" etc. Perhaps your textbook has a vocabulary section.

OpenStudy (skullpatrol):

Just say the number of units a given point on F(x) moves up and to the left.

OpenStudy (anonymous):

oh, ok. and i would look it up in a textbook, but its a little bit more complicated than that haha. i do online school which means my textbook in no way corresponds to my schoolwork nor have i ever used it so have no idea where my current unit would be in the book. thanks (:

OpenStudy (skullpatrol):

By a "given" point I mean choose a convenient point to describe :)

OpenStudy (skullpatrol):

The words you use to describe what is going on are not as important as you knowing what effect each different change in the original equation F(x) = 1/x

OpenStudy (anonymous):

wait, what do you mean? you mean like what the +4 means in the equation verses having nothing there? what does that mean lol

OpenStudy (anonymous):

@skullpatrol ?

OpenStudy (skullpatrol):

Yes, that is exactly what I mean.

OpenStudy (skullpatrol):

btw, thanks for the testimonial :)